OpenStudy (anonymous):

What are the zeroes of the function? What are their multiplicities? f (x) = 4x3 – 20x2 + 24x

OpenStudy (hba):

\[4x^3 – 20x^2 + 24x=4x(x^2-5x+6)\] is a start then factor the quadratic to find the other two zeros

OpenStudy (hba):

there will be 3 zeros, each with multiplicity 1

OpenStudy (anonymous):

The numbers –3, –2, and 0 are zeroes of multiplicity 1. The numbers 3, 2, and 0 are zeroes of multiplicity 2. The numbers –3, –2, and 0 are zeroes of multiplicity 2. The numbers 3, 2, and 0 are zeroes of multiplicity 1.

OpenStudy (anonymous):

so which one would it be ?

OpenStudy (hba):

yes. multiplicity means how many times a root is repeated. notice in 4x*(x^2-5x+6)= 0 the x=0 makes this 4*0*(stuff) which will be 0 x=0 makes (x^2-5x+6) = 6 so x=0 works as a root only once. so it is multiplicity 1. if you had x*x =0 then either copy of x could be 0, and it is multiplicity 2

OpenStudy (hba):

you can use the sum of roots = -b/a to check the remaining choices you match x^2-5x+6 to ax^2 +bx +c and see a=1, b= -5 so -b/a= -(-5)/1= +5 the sum of the roots (other than 0) add to 5

OpenStudy (anonymous):

so then it is either a or d...

OpenStudy (anonymous):

so it's d?

OpenStudy (hba):

or you could factor x^2-5x+6 the + in +6 means both factors have the same sign list the pairs of factors of 6 (1,6) (2,3) which pair add up to 5? 2,3 the sign of -5 (from -5x) says the biggest factor is negative, so both a - we get (x-2)(x-3)=0 solve for the roots x-2=0 or x=2 x-3=0 or x=+3

OpenStudy (anonymous):

thanks (: